# Characterizing Continual Learning Scenarios and Strategies for Audio Analysis

### v1

@@ Abstract
Audio analysis is useful in many application scenarios. The state-of-the-art audio analysis approaches assume that the data distribution at training and deployment time will be the same. However, due to various real-life environmental factors, the data may encounter drift in its distribution or can encounter new classes in the late future. Thus, a one-time trained model might not perform adequately. In this paper, we characterize continual learning (CL) approaches in audio analysis. In this paper, we characterize continual learning (CL) approaches, intended to tackle catastrophic forgetting arising due to drifts. As there is no CL dataset for audio analysis, we use DCASE 2020 to 2023 datasets to create various CL scenarios for audio-based monitoring tasks. We have investigated the following CL and non-CL approaches: EWC, LwF, SI, GEM, A-GEM, GDumb, Replay, Naive, cumulative, and joint training. The study is very beneficial for researchers and practitioners working in the area of audio analysis for developing adaptive models. We observed that Replay achieved better results than other methods in the DCASE challenge data. It achieved an accuracy of 70.12% for the domain incremental scenario and an accuracy of 96.98% for the class incremental scenario.
@@ I Introduction
Audio plays an important role in machine learning and deep learning applications, contributing to various tasks such as speech recognition [1], audio surveillance [2], sound classification, audio signal processing [3], music generation, and environmental sound analysis. In this paper, we characterize continual learning (CL) approaches for audio analysis with a case study of audio-based monitoring. Audio-based monitoring offers distinct advantages such as low-cost deployment, non-invasiveness, remote operation, and the ability to capture rich contextual information. With advancements in sensor technologies and the widespread availability of audio data, audio-based monitoring has become increasingly prevalent across various domains, including public surveillance, environmental sound monitoring, wildlife conservation, and healthcare.
In the real world, data cannot be obtained at once; it usually comes over time. We need to keep updating the models with new data. However, retraining with new data leads to a phenomenon known as catastrophic forgetting [4][5][6][7]. To mitigate this challenge, we need to use approaches that can remember everything we learned once. Continual learning (CL) is an approach that allows the model to learn new data and remember the previous learning simultaneously [8, 9]. Different CL approaches are used to handle catastrophic forgetting. The approaches compared in this paper are EWC [10], SI [11], LwF [12][13], GEM [14], A-GEM [15], GDumb [16], Replay, Naive, cumulative, and Joint Training. We have used ResNet50 as the backbone model for these approaches.
CL has three fundamental scenarios: task-incremental, domain-incremental, and class-incremental scenario. In this paper, the dataset is arranged in two criteria: domain incremental and class incremental. In domain incremental, the data consists of data samples in different batches or tasks, and each task has similar classes but a different distribution of data samples. While in class incremental, the tasks can have different classes associated with the data distribution. Different CL scenarios are created using data from DCASE [^1] https://dcase.community/ challenge, which provides a huge publicly available audio dataset with annotations.
@@
  • We have given a comparative study of different CL approaches, namely EWC, LwF, SI, GEM, A-GEM, GDumb, Replay, Naive, cumulative, and JointTraining with maximum possible variations of their parameter values to get the best performance of the particular approach for two CL scenarios like domain incremental and class incremental.
  • We have prepared a domain and class incremental scenarios dataset using the publicly available DCASE audio dataset of 2020 to 2023 challenges.
The rest of the paper is structured as follows. Section II contains the literature survey on monitoring and CL approaches, followed by section III, which gives the basics of CL methods and scenarios. Section IV consists of the basics of CL and Non-CL baselines. Section V includes a detailed description of the dataset used. Section VI consists of the experimental setup and results, followed by a discussion in section VII, and the conclusion in section VIII.
II Literature Review
This section provides a brief overview of CL approaches in the context of audio analysis. CL methods enable machine learning models to learn sequentially from a continuous stream of data. CL allows them to acquire new skills and accumulate knowledge over time without the risk of forgetting previously learned experiences, commonly known as catastrophic forgetting [17]. In the audio domain, continual learning becomes pertinent due to the sequential, dynamic, and evolving nature of sound environments, which requires adaptive models. By continuously updating models with new audio samples, CL algorithms can learn to recognize and classify new audio events, adapt to new acoustic conditions, and generalize across different audio contexts. The audio context or applications seen so far are sound classification [18], audio analysis/classification [19] [20][21][22][23], sound event detection [24][25], audio classification [26], audio captioning [27], audio-visual learning [28][29][30], fake audio detection [31], speech recognition [32], etc.
A summary of CL works in the audio domain is provided in Table I. From Table I, we observe that most of the audio-based CL applications focus on audio scene classification, with very few addressing anomaly detection. It can be inferred that the DCASE anomaly dataset utilized in this paper has not been previously employed by others. Although some studies have utilized the DCASE dataset, they have done so for different applications, such as audio scene classification, rather than anomaly detection. For instance, in [33], the DCASE challenge anomaly dataset was utilized for anomaly detection, although in a different manner. They treated subsequent machine types as the drift, whereas we considered naturally occurring drift. Further, we can observe that most works focused on handling the unseen classes or working in class incremental scenarios, and very few handle the domain shift. Moreover, almost all works focus on either scenario, i.e., domain or class incremental scenario, but not both. In this paper, we have studied both scenarios and performed a comparative analysis of various state-of-the-art continual learning strategies.
@@ II Literature Review
Wang et. al.[36] Architectural CI
FSD-MIX-CLIPS (1s),
FSD-MIX-SED (10s)
Audio scene classification
@@ III Continual learning
Continual learning (CL) is a rapidly emerging research direction to mitigate catastrophic forgetting and mimic the human way of learning [8]. It refers to the learning in which a machine continuously learns from the incoming data. Unlike traditional deep learning, where the complete data is accessible for training at the start, in CL, the data comes sequentially, also known as tasks or episodes. CL allows the acquisition of new knowledge while retaining the existing knowledge. However, there is a tradeoff between how well a model can retain past knowledge and learn new things, called the stability-plasticity dilemma [49]. The plasticity shows the ability of the model to learn new tasks, and the stability refers to maintaining the performance of previously learned tasks upon learning new concepts [50]. Efforts to increase either of them hinder the other and vice versa [9]. More favor for plasticity or stability also depends on the targeted application. There are various CL strategies to mitigate forgetting and allow knowledge accumulation. We discuss them in the following sections. †Section III-A and Section III-B.
III-A Continual learning scenarios
In analysis, particularly audio-based monitoring using anomaly detection, two situations are most prevalent: new samples of the existing class due to various shifts and new intruder/event classes. The first type deals with data with shifted distribution and is termed as ‘new instance’ (NI) scenario; the other type deals with new classes in subsequent episodes and is termed as ‘new class’ (NC) scenario in CL research. Allowing various permutations and combinations on NI, NC, and unique tasks, various CL scenarios are defined across multiple research domains [51], as follows:
III-A1 Data-incremental scenario
Data-incremental scenario aims to continuously learn from new streams of data from an identical underlying distribution during sequential model training. It does not consider new classes or considerable shifts of data distribution and is thus expected to be the easiest among other CL scenarios. It is common for agriculture experts to generate annotations for a dataset in multiple batches, and thus each batch can be considered as a sample of the same data distribution.
III-A2 Domain-incremental scenario
@@
Domain-incremental (DI) scenario deals with situations where the task remains the same, but the context or the information presented changes over time due to the non-stationary environment. In other words, it is an NI scenario with a single task. The data distribution changes, but the classes remain the same for all subsequent episodes. Domain-incremental learning aims to continuously update the model with new streams of data from shifted distributions. Examples of various shift sources in monitoring include variations due to weather, seasons, geo-location, surrounding noise levels across the day, etc.
III-A3 Class-incremental scenario
@@
Class-incremental (CI) scenario is an NC scenario with a single task over the subsequent episodes. It deals with continuously extending the model to unseen classes with new streams of data. Usually, the classes in subsequent episodes are disjoint, and hence the knowledge to learn is also non-overlapping, causing this scenario to be the hardest among CL scenarios. An example of class-incremental learning in real-world monitoring is to identify a cat and dog in one episode and then a cow and horse in the other episode [52].
III-A4 Task-incremental scenario
Task-incremental (TI) scenario refers to learning different kinds of tasks over the episodes. The tasks may also be highly uncorrelated. Each task observed as an episode, usually has a disjoint label space. A real-world example of task-incremental learning is the sequential learning of different sports or musical instruments over time (episode). The information regarding which sports or instrument is the current task can be provided (task-aware) or not required (task-agnostic). The CL literature is ambiguous for task-incremental settings as class-incremental and domain-incremental learning are frequently referred to as task-incremental learning by different research communities [52].
III-A5 Hybrid scenarios
In the real world, it is highly unrealistic that the model will always encounter one type of incremental scenario. It is quite possible that the model can observe new classes as well as domain shifts. The model developed for handling class-incremental situations may not perform well for domain-incremental situations. Thus, there are attempts in the literature to evaluate the same model on different incremental scenarios [53]; however, a separate evaluation is followed for different kinds of incremental learning. In contrast, episodes with a mix of incremental scenarios are closer to real-life situations. Thus it becomes important to also evaluate the CL model in hybrid-incremental settings, i.e., a sequence of episodes with interleaved class-incremental and domain-incremental can be considered for evaluation.
Figure 1: Overview of a continual learning pipeline
III-B Continual learning techniques
@@
Existing CL techniques to handle catastrophic forgetting can be broadly classified into three categories, viz., rehearsal, regularization, and architectural. In rehearsal-based techniques, a small memory is used to store previous episode data and then replayed along with current data to overcome forgetting. The previous data can be stored and replayed in various forms, including raw samples, deep feature descriptors, and generative models to generate samples or features. In contrast, Regularization-based techniques avoid storing past samples and add a regularization term in the loss function or regularize the learning rate to control drastic deviations in learned weights to minimize forgetting. Further, data-focused and prior-focused regularization have been explored. Data-focused approaches use knowledge distillation to transfer past learning to the current model, whereas the prior-focused methods define the importance of the network’s parameters, which is then used to penalize the shift of important parameters [10]. Architectural techniques mainly aim to keep some network parameters reserved for each episode. Two design choices are explored in this category: fixed network capacity and dynamic network capacity. Lastly, hybrid approaches, i.e., a combination of two or more of the above categories, are also popularly explored in various applications [53].
IV CL and non-CL Baselines
The CL research community has contributed various techniques in the above-mentioned categories. Their usefulness has been analyzed in various applications, including object detection [54], robotics [55], medical image analysis [56, 51], etc. For the current application of audio analysis, we have evaluated the following CL and non-CL baselines.
IV-A Continual learning baselines
IV-A1 Elastic Weight Consolidation
@@
IV-A2 Learning without Forgetting
Learning without Forgetting (LwF) is a regularization method aimed at addressing the challenge of forgetting in neural networks by enforcing stability in the output or predictions. Notably, LwF prioritizes training efficiency, demonstrating faster results compared to the joint training approach. This is a CL approach, specifically designed to utilize only new data, assuming the unavailability of past data used for initial network pre-training. This sets LwF apart from other CL methods, which typically focus on leveraging prior knowledge to facilitate the learning of new tasks [12][13].
IV-A3 Synaptic Intelligence
@@
IV-A4 Gradient Episodic Memory
@@
IV-A5 Averaged Gradient Episodic Memory
Averaged Gradient Episodic Memory (A-GEM) tries to alleviate the computational cost caused by GEM. A-GEM can be defined as the more efficient version of GEM. Unlike GEM, which minimizes the loss of each previous tasks at each training step by adjusting the gradient of each current task, A-GEM minimizes the average episodic memory loss over the previous tasks at each training step [15].
IV-A6 Greedy sampler and Dumb learner
In Greedy sampler and Dumb learner (GDumb), The sampler greedily stores the part of data samples as they are encountered while ensuring a balanced representation of classes, and the learner which is a neural network trains it dumbly using all the samples available in the memory. Therefore it is called GDumb. Although GDumb is not meant to handle the continual learning scenarios, it gives convincing results in comparison to the CL strategies [16].
IV-A7 Replay
The replay strategy can use different types of replay, like, experience replay where the old stored samples from the memory are used for training with the new samples, generative replay where the old data is generated using some generative models and trained with the new encountered data samples etc. Here we have used simple replay.
IV-B Non-Continual learning baselines

### v2

@@ Abstract
Audio analysis is useful in many application scenarios. The state-of-the-art audio analysis approaches assume the data distribution at training and deployment time will be the same. However, due to various real-life challenges, the data may encounter drift in its distribution or can encounter new classes in the late future. Thus, a one-time trained model might not perform adequately. Continual learning (CL) approaches are devised to handle such changes in data distribution. There have been a few attempts to use CL approaches for audio analysis. Yet, there is a lack of a systematic evaluation framework. In this paper, we create a comprehensive CL dataset and characterize CL approaches for audio-based monitoring tasks. We have investigated the following CL and non-CL approaches: EWC, LwF, SI, GEM, A-GEM, GDumb, Replay, Naive, Cumulative, and Joint training. The study is very beneficial for researchers and practitioners working in the area of audio analysis for developing adaptive models. We observed that Replay achieved better results than other methods in the DCASE challenge data. It achieved an accuracy of 70.12% for the domain incremental scenario and an accuracy of 96.98% for the class incremental scenario.
@@ I Introduction
Audio plays an important role in machine learning and deep learning applications, contributing to various tasks such as event detection [1], speech recognition [2], audio surveillance [3], sound classification [4], audio signal processing [5], music generation, and environmental sound analysis. In this paper, we characterize continual learning (CL) approaches for audio analysis with a case study of audio-based monitoring. Audio-based monitoring offers distinct advantages such as low-cost deployment, non-invasiveness, remote operation, and the ability to capture rich contextual information [6, 7, 8]. With advancements in sensor technologies and the widespread availability of audio data, audio-based monitoring has become increasingly prevalent across various domains, including public surveillance, environmental sound monitoring, wildlife conservation, and healthcare.
State-of-the-art audio analysis models follow one-time training with all the available data and employ the trained model for all future inferences. The assumption of prior availability of all the data to train a deep model is not practical. Data usually comes over time in real-world environments, and the model needs to be updated with the new batch of data [9, 10]. However, as the deep models learn new concepts from newly available data, they tend to forget the past learning and hence perform poorly on past data, leading to a phenomenon known as catastrophic forgetting [11, 12, 13, 14]. To mitigate the effect of catastrophic forgetting, a human-like lifelong learning paradigm, popularly termed as Continual Learning (CL) is gaining major attention recently. CL offers to acquire new skills and accumulate knowledge over time without the risk of forgetting previously learned experiences. Consequently, several CL strategies have been proposed in various research fields, including object detection [15], robotics [16], medical image analysis [17, 18], etc.
Especially in the audio domain, CL becomes pertinent due to the streaming and dynamically evolving nature of sound environments. By continuously updating the deployed model with new audio samples, a CL strategy can be leveraged to recognize and classify new audio events, adapt to new acoustic conditions, and generalize across different audio contexts. CL has been exploited in audio applications including sound classification [4], audio analysis/classification [19, 20, 21, 22, 23], sound event detection [24, 25], audio captioning [26], audio-visual learning [27, 28, 29], fake audio detection [30], speech recognition [31], etc. Nevertheless, efforts to adopt CL strategies in the audio monitoring application are still in their early stage. The authors mostly use custom datasets in ad-hoc scenarios to demonstrate their performance. There is a need for a systematic evaluation framework and scenario-rich dataset to assess the current state of CL works in the field of audio analysis.
We curate a sequential dataset in “domain-shift” as well as “new-class” situations with different CL scenarios using data from DCASE challenges [^1] https://dcase.community/. DCASE is a huge publicly available audio dataset repository with annotations. Although Zhou et al. [32] utilize anomaly dataset from the DCASE2020 task2 challenge, they consider drift as a change of machine type in the anomaly detection application, whereas we considered naturally occurring drifts such as weather conditions, machine load over time, etc. Further, we identify and benchmark popular CL strategies for audio-based monitoring application including EWC [33], SI [34], LwF [35][36], GEM [37], A-GEM [38], GDumb [39], Replay[40], Naive, Cumulative, and Joint training. We conduct a thorough analysis of the performance and examine how well catastrophic forgetting is controlled by these CL strategies in different dynamically changing audio environment settings.
@@
  • We have given a comparative study of different CL approaches, namely EWC, LwF, SI, GEM, A-GEM, GDumb, Replay, Naive, Cumulative, and Joint training for audio-based monitoring tasks.
  • We have prepared a domain and class incremental scenarios dataset using the publicly available DCASE challenge datasets from the years 2020 to 2023.
The rest of the paper is structured as follows. Section II gives the details of CL methods and scenarios. Section III consists of the basics of CL and Non-CL baselines. Section IV includes a detailed description of the dataset used. Section V consists of the experimental setup and results, followed by a discussion in section VI, and the conclusion in section VII.
@@
Multi-label audio classification
Wang et. al.[45] Architectural CI
FSD-MIX-CLIPS (1s),
FSD-MIX-SED (10s)
@@ II Continual learning
CL is a rapidly emerging research direction to mitigate catastrophic forgetting and mimic the human way of learning in deep neural networks [47]. It refers to the learning strategy in which a machine continuously accumulates knowledge from the incoming data. Unlike traditional deep learning, where the complete data is accessible for training at the start, in CL, the data comes sequentially, also known as tasks or episodes. CL allows the acquisition of new knowledge while retaining the existing knowledge. However, there is a tradeoff between how well a model can retain past knowledge and learn new things. It is called the stability-plasticity dilemma [48]. The plasticity shows the ability of the model to learn new tasks, and the stability refers to maintaining the performance of previously learned tasks upon learning new concepts [49]. Efforts to increase either of them hinder the other, and vice versa [50]. More favor for either plasticity or stability also depends on the targeted application.
Various CL strategies have been formulated in literature to mitigate forgetting and facilitate knowledge accumulation. Fig. 1 details broader modules of a CL framework, including CL scenario, CL data stream, CL strategy, evaluation scheme, and finally, available baselines. Depending upon what kind of change is expected, i.e., new class, new instances, shifted distribution, entirely new task, etc., there can be various settings, also known as CL scenarios. Once the possible change in data is identified for the application, a dataset stream containing multiple tasks, each having a train and test split, is prepared. Then, a suitable CL strategy from various categories is employed in the deep model to refrain from a drop in performance on past episodes upon learning new episodes. Lastly, the effectiveness of a particular CL strategy is analyzed against other CL and non-CL strategies with the help of specifically designed metrics to measure forgetting and knowledge transfer. We discuss major CL scenarios and strategies in sections II-A and II-B, respectively.
II-A Continual learning scenarios
In continual learning, a “scenario” is the specific way in which new tasks are presented to the model over time. It describes the conditions under which the model learns and adapts, simulating how learning happens in the real world, where new information comes in a sequence. In analysis, particularly audio-based monitoring using anomaly detection, two situations are most prevalent: new samples of the existing class with various shifts and new intruder/event classes. Based on these situations, various CL scenarios are defined across multiple research domains [18], out of which two majorly important and widely used scenarios are discussed below.
II-A1 Domain-incremental scenario
@@
Domain-incremental (DI) scenario deals with situations where the job remains the same, but the context or the information presented changes over time due to the non-stationary environment. The data distribution changes, but the classes remain the same for all subsequent tasks. The domain-incremental scenario [51] allows us to continuously update the model with new streams of data from shifted distributions. Examples of various shift sources in monitoring include variations due to weather, seasons, geo-location, surrounding noise levels across the day, etc.
TABLE II: Summary of major CL scenarios
CL Scenario Summary
DI
Each task contains new data distribution with the
same classes per task, task ID is not provided
CI
Each task contains a new set of classes per task,
task ID is inferred
II-A2 Class-incremental scenario
@@
Class-incremental (CI) scenario [52] is a scenario with a single job over the subsequent tasks. It deals with continuously extending the model to unseen classes with new streams of data. Usually, the classes in subsequent tasks are disjoint; hence, the knowledge to learn is also non-overlapping, causing this scenario to be the hardest among CL scenarios. An example of class-incremental learning in real-world monitoring is to identify a cat and dog in one episode, and then a cow and horse in the next episode [53].
There are other CL scenarios available in the literature like Data-incremental, Task-incremental, and Hybrid scenarios. The data-incremental scenario allows us to continuously learn from new streams of data from an identical underlying distribution during sequential model training. It does not consider new classes or considerable shifts of data distribution and thus can also be considered as a domain-incremental scenario where the data distribution remains unchanged. The task-incremental (TI) scenario refers to learning different kinds of jobs over subsequent tasks. The jobs may also be highly uncorrelated. Each job is observed as a task, and it usually has a disjoint label space. The task-incremental scenario can be considered the class-incremental scenario where the subsequent tasks are observed as dissimilar jobs and classes associated with the tasks are disjoint. It can also be considered the domain-incremental scenario where the subsequent tasks are observed as similar jobs, and classes associated with them are unchanged. A real-world example of task-incremental learning is the sequential learning of different sports or musical instruments over time. In the real world, it is highly unrealistic that the model will always encounter one type of incremental scenario. It is quite possible that the model can observe new classes as well as domain shifts. The model developed for handling class-incremental situations may not perform well for domain-incremental situations. Thus, there are attempts in the literature to evaluate the same model on a mix of incremental scenarios called hybrid scenarios [54]. The two major CL scenarios are summarized in table II.
Figure 1: Overview of a continual learning pipeline. This figure shows the available possible CL scenarios, the data preparation for different scenarios in subsets called tasks, broad categories of CL strategies based on literature, the well-known metrics used to evaluate the performances, and different ways of comparative analysis.
II-B Continual learning techniques
@@
Existing CL techniques to handle catastrophic forgetting can be broadly classified into three categories [55][56], viz., rehearsal [37][38][40][57], regularization [33][34][35][36], and architectural [58][59]. In rehearsal-based techniques, a small memory is used to store previous task data and then replayed along with current data to overcome forgetting. The previous data can be stored and replayed in various forms, including raw samples, deep feature descriptors, and generative models to generate samples or features. In contrast, Regularization-based techniques avoid storing past samples and add a regularization term in the loss function or regularize the learning rate to control drastic deviations in learned weights to minimize forgetting. Further, data-focused and prior-focused regularization have been explored. Data-focused approaches use knowledge distillation to transfer past learning to the current model, whereas the prior-focused methods define the importance of the network’s parameters, which is then used to penalize larger deviation in important parameters [33]. Architectural techniques mainly aim to keep some network parameters reserved for each episode. Two design choices are explored in this category: fixed network capacity and dynamic network capacity. Lastly, hybrid approaches, i.e., a combination of two or more of the above categories, are also popularly explored in various applications [54].
II-C Continual Learning for Audio Analysis
Table I summarises various studies that apply CL to audio applications and utilize CL strategies, scenarios, and datasets. For audio scene classification, researchers have employed CI and TI scenarios to handle the dynamic nature of audio data. The works [4][25][41][42][43] and [44] focused on CI scenario. In contrast, Karam et al. [20] targeted the TI scenario for classification. In multi-label audio classification [19][45][23], researchers have used CI scenario on AudioSet, ESC50, and FSD-MIX-CLIPS datasets. Authors used CI scenarios for multi-modal classification, which involves audio and video data [28] [29]. They used the VGGSound and AVE-CI datasets to ensure the model could learn from both audio and visual inputs without degrading performance. Kim et al. [46] applied the TI scenario on diverse datasets such as VGGSound, MACS, and ClothoV2 datasets. Berg et al. [26] used Clotho and AudioCaps datasets to maintain accuracy in generating captions for audio clips. Ma et al. [30] focused on the ASVSpoof2019 dataset for fake audio detection, where drift is considered as the different types of spoofing attacks with major differences. Zhou et al. [32] used the DI scenario for anomaly sound detection in the DCASE 2020 challenge, addressing the challenge of subsequent machine-type drift and ensuring robust anomaly identification over time.
From the summary provided in Table I, we can observe that most of the works focus on audio scene classification. Further, most “new class occurrence” type situation (class-incremental) is considered in these works, whereas the “drifted data” situation (domain-incremental), which is a more prevalent and critical issue in outdoor audio scene analysis, is largely ignored. Less exploration is also attributed to the lack of suitable public datasets for the evaluation of a CL framework.
III Baselines
The CL research community has contributed various CL techniques in the categories mentioned in Section II-B. Their effectiveness is also evaluated against non-CL approaches that try to mitigate forgetting using retraining or exploiting all the datasets together. However, these approaches do not fit with human-like learning but may perform well at the cost of large storage and computation facilities. For the current application of audio analysis, we have evaluated the following CL and non-CL baselines.
III-A Continual learning baselines
This paper targets a few major and frequently used CL baselines to conduct a comparative study on the dataset. These baselines are briefly discussed below.
III-A1 Elastic Weight Consolidation
@@
III-A2 Learning without Forgetting
Learning without Forgetting (LwF) is a regularization method aimed at addressing the challenge of forgetting in neural networks by enforcing stability in the output or predictions. Notably, LwF prioritizes training efficiency, demonstrating faster results compared to the joint training approach. This is a CL approach specifically designed to utilize only new data, assuming the unavailability of past data used for initial network pre-training. This sets LwF apart from other CL methods, which typically focus on leveraging prior knowledge to facilitate the learning of new tasks [35][36].
III-A3 Synaptic Intelligence
@@
III-A4 Gradient Episodic Memory
@@
III-A5 Averaged Gradient Episodic Memory
Averaged Gradient Episodic Memory (A-GEM) tries to alleviate the computational cost caused by GEM. A-GEM can be defined as the more efficient version of GEM. Unlike GEM, which minimizes the loss of each previous task at each training step by adjusting the gradient of each current task, A-GEM minimizes the average episodic memory loss over the previous tasks at each training step [38].
III-A6 Greedy sampler and Dumb learner
In Greedy sampler and Dumb learner (GDumb), the sampler greedily stores the part of data samples as they are encountered while ensuring a balanced representation of classes, and the learner, which is a neural network, trains it dumbly using all the samples available in the memory. Therefore it is called GDumb. Although GDumb is not meant to handle the continual learning scenarios, it gives convincing results in comparison to the CL strategies [39].
III-A7 Replay
The replay strategy can use different types of replay, like, experience replay [40] where the old stored samples from the memory are used for training with the new samples, generative replay [61] where the old data is generated using some generative models and trained with the new encountered data samples etc.
III-B Non-Continual learning baselines